OpenStudy (anonymous):

What is the area of a triangle with side lengths 5,12 and 13?

OpenStudy (yrelhan4):

You should realise this is an equilateral triangle with hypotensue 13. the height/base as 5/12. 1/2 *12*5 should give you the area.

OpenStudy (gorv):

it is a right angle triangle so area =1/2*base*height =1/2*12*5

OpenStudy (gorv):

=30